# Precise Interception Flight Targets by Image-based Visual Servoing of Multicopter

### v1

@@ Abstract
Interception of low-altitude intruding targets with low-cost drones equipped strapdown camera presents a competitive option. However, the malicious maneuvers by the non-cooperative target and the coupling of the camera make the task challenging. To solve this problem, an Image-Based Visual Servoing (IBVS) control algorithm based on proportional navigation guidance with field-of-view holding capability is designed. The proposed controller reduces the miss distance while improving the stability of the visual servo system during interception. Software-in-the-loop (SITL) simulation experiments show a 72.8% reduction in the circular error probability (CEP) compared to the most recent study. This improvement enhances interception accuracy from the decimeter to the centimeter level. Real-world experiments further validate the effectiveness of the proposed algorithm.
@@ I Introduction
THE presence of non-cooperative targets at low altitudes poses a significant threat to flight safety and hinders the development of low-altitude economies[1, 2, 3, 4]. Commonly used countermeasures such as radio-frequency (RF) signal jamming[5, 6] and high-energy weapon shoot-downs[7, 8] are effective but have detrimental environmental impacts. The capture methods[9, 10] are environmentally friendly but less effective against moving targets. All these methods exhibit notable deficiencies when dealing with moving targets. The use of sensor-equipped unmanned aerial vehicles (UAVs) to intercept intruding targets has gained significant attention due to their rapid deployment, safety, and cost-effectiveness. Camera-based solutions, in particular, have shown great potential due to their low cost, lightweight, and high versatility[11, 12, 13].
@@
In recent years, much research has focused on camera-equipped multicopters for intercepting non-cooperative targets. These methods are primarily divided into two categories: Location estimation-based and IBVS-basd. Common Location estimation-based methods employ binocular cameras[11, 14], the Kalman filter[13, 15], and geometry-based approaches[16] for target estimation. After that, a trajectory is planned for interception. However, these methods are error-sensitive and limited by sensor perception range and computational performance. Yet the IBVS methods have a simple structure that calculates the control quantity directly from the sensing module. So it offers advantages such as insensitivity to modeling and calibration, fast response, and extended sensing capabilities. These advantages make it a highly competitive solution for interception tasks.
@@
Multicopter IBVS faces significant challenges in precise interception of non-cooperative flight targets. During interception, the observable field of view (FOV) is susceptible to the multicopter’s attitude, making it challenging to maintain the 2D visibility required for IBVS [17]. Interception accuracy is further affected by target maneuvers, image processing delay, multicopter dynamics delay, and guidance strategy. In our previous work[18, 19], four and six degree of freedom (DOF) IBVS controllers were designed for low and high speed target interception, respectively. Delayed Kalman filtering (DKF) was proposed in [19] for mitigating image processing delay. Some other studies focus on target observation innovations. A virtual-plane IBVS scheme [20] decouples multicopter dynamics and uses an improved image error term to reduce the impact of multicopter rotation during 3D target chasing. A pseudo-linear Kalman filter [21] and a 3D helical guidance law enhances observability for intruding multicopter tracking. Classic guidance laws, such as the pursuit guidance (PG) method [22], aim for lower off-target interception but are less efficient for moving targets. These methods can cause multicopter overload saturation or loss of the target due to image processing latency, inappropriate guidance strategies, and target maneuvers. No related work has focused on the bad impact of overload on interception accuracy. An ideal flat trajectory at the end phase of interception reduces direction angle measurement errors from image processing latency and provides a greater dynamics margin to handle target maneuvers.
@@
In this paper, an IBVS controller is designed for precision interception of non-cooperative flight targets, inspired by bird-catching behaviors. Birds rely solely on vision for efficient hunting, similar to IBVS methods for multicopter interception tasks (see Fig. 1). Specifically, the trajectory of a peregrine falcon hunting mallard duck could described by proportional navigation guidance (PNG) [23], inspired our use of PNG for our tasks. PNG, widely used in missile guidance for decades, flattens the end-phase trajectory and reduces the miss distance[24]. Additionally, a FOV holding controller mimics the bird’s neck rotation for prey tracking. So the proposed controller can continuously and stably track the target and precise interception like a peregrine falcon. Overall, the contributions of this paper are summarized as follows:
@@
    The proposed algorithm achieved a Circular Error Probability (CEP) of 0.089 meters in typical tests, improving interception accuracy from the decimeter level of recent algorithms [19] to the centimeter level. CEP is a measure where there is a 50% probability that the interceptor will hit within a circle centered on the target with the given radius [25].
@@
    A FOV holding controller that incorporates PNG and multicopter dynamics is designed to enhance robustness in target tracking.
@@
    Real and simulated experiments, with extensive comparisons across various scenarios, validate the effectiveness and accuracy of the proposed algorithms.
@@ II-A Coordinate systems and relative motion models
There are four coordinate systems used, as shown in Figure 2: the Earth-Fixed Coordinate System (EFCS): $o_{e}-x_{\mathrm{e}}y_{e}z_{\mathrm{e}}$, the Body Coordinate System (BCS): $o_{b}-x_{b}y_{b}z_{b}$, Camera Coordinate System (CCS): $o_{c}-x_{c}y_{c}z_{c}$ and Image pixel coordinate system (IPCS): $o_{i}-uv$.
@@ II-B Multicopter Dynamics Model
where ${}^{\mathrm{e}}\mathbf{p}_{\mathrm{M}}$ and ${}^{\mathrm{e}}\mathbf{v}_{\mathrm{M}}$ is the multicopter’s position and velocity in the EFCS, respectively. $\mathbf{g}$ denotes the acceleration due to gravity, typically represented as $\left[\begin{array}[]{ccc}0&0&g\end{array}\right]^{\mathrm{T}}$ with $g\approx$ 9.8 $m/s^{2}$. ${}^{\mathrm{e}}\mathbf{f}\in\mathbf{R}^{3}$ is the controlled lift of the multicopter, directed opposite to the z-axis of the airframe. $\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\in\mathrm{SO}\left(3\right)$ represents the rotation from the BCS to the EFCS; ${}^{\mathrm{b}}\omega$ is the angular velocity in the BCS, and $\left[{}^{\mathrm{b}}\bm{\omega}\right]_{\times}$ is the cross-product matrix associated with ${}^{\mathrm{b}}\bm{\omega}$, where $\left[\cdot\right]_{\times}$ denotes the matrix such that $\left[\mathbf{x}\right]_{\times}\mathbf{y}=\mathbf{x}\times\mathbf{y}$ for any $\mathbf{x},\mathbf{y}\in\mathbf{R}^{3}$. $\mathbf{J}$ is the moment of inertia of the multicopter, $\mathbf{G}_{\mathrm{a}}$ represents the gyroscopic moment, and ${}^{\mathrm{b}}\mathbf{M}_{\mathrm{d}}$ denotes the aerodynamic moment associated with propeller steering.
@@ II-C Camera Image Model
where ${}^{\mathrm{c}}\tilde{\mathbf{v}}=\left[{}^{\mathrm{c}}\mathbf{v}^{\mathrm{T}}\,\,{}^{\mathrm{c}}\mathbf{\omega}^{\mathrm{T}}\right]^{\mathrm{T}}$ reprents the instantaneous linear velocity angular velocity of the camera. ${\bar{\mathbf{e}}}=\left[\bar{e}_{x}\,\,\bar{e}_{y}\right]^{\mathrm{T}}=\left[e_{x}/f_{oc}\,\,e_{y}/f_{oc}\right]^{\mathrm{T}}$ represents the normalized image error of features in IPCS, and $f_{\mathrm{oc}}$ is the focal length of the camera.
@@ II-D Framework overview
The system framework is shown in Fig. 3. First, the image is processed using a target detection algorithm [28], and corrections are estimated using the DKF [19]. The IBVS controller algorithm then calculates the desired lift $f_{\mathrm{d}}$ and angular velocity $\omega_{\mathrm{d}}$ for the multicopter flight controller. Finally, the controller generates PWM signals for the motors to execute the target interception.
@@ III-A Guidance Algorithm Design
$q_{z}=\arctan\left(\frac{n_{ty}}{n_{tx}}\right);q_{z}\in(-\pi,\pi).$
And normalized velocity direction vector $\mathbf{n}_{\mathbf{v}}=\left[n_{vx}\,\,n_{vy}\,\,n_{vz}\right]^{\mathrm{T}}$. Thus, $\sigma_{y}$ and $\sigma_{z}$ are the velocity angles in the vertical and horizontal planes which are donated as:
@@
$\sigma_{z}=\arctan\left(\frac{n_{vy}}{n_{vx}}\right);\sigma_{z}\in(-\pi,\pi).$
@@
$\left\{\begin{array}[]{c}\sigma_{\mathrm{yd}}=K_{y}(q_{y(k)}-q_{y(k-1)})+\sigma_{y(k-1)}\\ \sigma_{\mathrm{zd}}=K_{z}(q_{z(k)}-q_{x(k-1)})+\sigma_{z(k-1)}\\ \end{array}\right.$ (9)
where $K_{y}$ and $K_{z}$ are proportionality coefficient for the vertical and horizontal planes. Timothy J. Layman et al. shows that a value of about 3 is appropriate when multirotor pursuit[29].
@@
where $\mathbf{n}_{\mathrm{vd}}=\left[\sin\sigma_{\mathrm{yd}}\cos\sigma_{\mathrm{zd}}\,\,\sin\sigma_{\mathrm{yd}}\sin\sigma_{\mathrm{zd}}\,\,\cos\sigma_{\mathrm{zd}}\right]^{\mathrm{T}}$. So the desired velocity can be get by designing the desired velocity magnitude.
@@ III-B Field of View Holding Controller
where $k_{\mathrm{p}}$ and $k_{\mathrm{d}}$ are small negative value. This controller improves the settling time and reduces the overshoot. Furthermore, it helps $e_{x}\rightarrow 0$ and is insensitive to pixel noise.
Changes in roll angle at greater distances from the target have less impact on the target in the FOV. Therefore, primary attention is given to changes in the pitch angle. To minimize the adverse effects of attitude changes, an acceleration controller is designed as follows:
@@
$\varDelta e_{y}\leqslant\frac{v_{0}\tan\left(\varDelta q_{\mathrm{y}}\right)}{\tan\left(\small{\frac{1}{2}\alpha_{\mathrm{vfov}}}\right)}\leqslant\varepsilon$ (16)
@@ III-C Attitude Loop Controller
$\mathbf{a}_{\mathrm{d}}=\mathbf{V}_{\mathrm{d}}-\mathbf{V}_{\mathrm{now}}$ (17)
@@
$L=\mathrm{tr}\left(\mathbf{I}-\mathbf{R}_{\mathrm{d}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right).$ (20)
Given that $\left\|\mathbf{I}-\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right\|_{\mathrm{F}}=\sqrt{2\mathrm{tr}\left(\mathbf{I}-\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right)}$, $L=\left\|\mathbf{I}-\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right\|_{\mathrm{F}}^{2}/2\geq 0$, the derivative of the candidate function is:
$\dot{L}$ $=-\mathrm{tr}\left(\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}[^{\mathrm{b}}\bm{\omega}]_{\times}\right)$ (21)
$=\mathrm{vex}\left(\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}-\mathbf{R}_{\mathrm{b}}^{\mathrm{eT}}\mathbf{R}_{\mathrm{d}}\right)^{\mathrm{T}}{\mathrm{b}}\bm{\omega}.$
To make $\dot{L}\leqslant 0$ for stability, the pitch and roll angular velocity controllers are designed as:

### v2

@@ Abstract
Vision-based interception using multicopters equipped strapdown camera is challenging due to camera-motion coupling and evasive targets. This paper proposes a method integrating Image-Based Visual Servoing (IBVS) with proportional navigation guidance (PNG), reducing the multicopter’s overload in the final interception phase. It combines smoother trajectories from the IBVS controller with high-frequency target 2D position estimation via a delayed Kalman filter (DKF) to minimize the impact of image processing delays on accuracy. In addition, a field-of-view (FOV) holding controller is designed for stability of the visual servo system. Experimental results show a circular error probability (CEP) of 0.089 m (72.8% lower than the latest relevant IBVS work) in simulations and over 80% interception success under wind conditions below 4 m/s in real world. These results demonstrate the system’s potential for precise low-altitude interception of non-cooperative targets.
@@ I Introduction
THE presence of non-cooperative targets at low altitudes poses a significant threat to flight safety and hinders the development of low-altitude economies[1, 2, 3, 4, 5]. Commonly used countermeasures such as radio-frequency (RF) signal jamming[6, 7] and high-energy weapon shoot-downs[8, 9] are effective but have detrimental environmental impacts including signal pollution in the airspace and fires on the ground[10, 11]. On the other hand, capture methods[12, 13, 14] are environmentally friendly but less effective against moving targets. All these methods exhibit notable deficiencies when dealing with moving targets. The use of sensor-equipped unmanned aerial vehicles (UAVs) to intercept intruding targets has gained significant attention due to their rapid deployment, safety, and cost-effectiveness. Camera-based solutions, in particular, have shown great potential due to their low cost, lightweight, and high versatility[15, 16, 17].
@@
In recent years, much research has focused on camera-equipped multicopters for intercepting non-cooperative targets. These methods are primarily divided into two categories: location estimation-based and IBVS-based. Common location estimation-based methods employ binocular cameras[15, 18], the Kalman filter[17, 19], and geometry-based approaches[20] for target estimation. Subsequently, a trajectory for interception is planned. However, these methods are error-sensitive and limited by the sensor perception range and computational performance. Yet the IBVS methods have a simple structure that calculates the control quantity directly from the sensing module. Therefore it offers advantages such as insensitivity to modeling and calibration, fast response, and extended sensing capabilities. These advantages make it a highly competitive solution for interception tasks.
@@
IBVS control of multicopter has been extensively studied in recent years and gradually there is some work for interception scenarios. Earlier methods support multicopter control such as stationary flight[21], autonomous landing[22] and translational rotation[23] by extracting image features. Recent IBVS-based multicopter approaches tend to use only the target’s bounding-box center point for airborne target tracking[24, 25] and interception[26, 27, 28]. Some studies focus on innovations in target observation during tracking. A virtual-plane IBVS scheme [25] decouples multicopter dynamics and uses an improved image error term to reduce the impact of multicopter rotation during 3D target chasing. A pseudo-linear Kalman filter [24] and a 3D helical guidance law enhance observability for intruding multicopter tracking. In our previous work[26, 28], four and six degree of freedom (DOF) IBVS controllers were designed for low and high speed target interception, respectively. Among them delayed Kalman filter (DKF) was proposed in [28] to mitigate the delay in image processing. While some classic guidance laws, such as the pursuit guidance (PG) method [27], aim at lower off-target interception but are less efficient for moving targets. These methods can cause multicopter overload saturation or loss of the target due to image processing latency, inappropriate guidance strategies, and target maneuvers. No related work has focused on the adverse impact of overload on interception accuracy.
Multicopter IBVS faces significant challenges in precise interception of non-cooperative flight targets. During interception, the observable field of view (FOV) is susceptible to the multicopter’s attitude, making it challenging to maintain the 2D visibility required for IBVS [29]. Interception accuracy is further affected by target maneuvers, image processing delay, multicopter dynamics delay, and guidance strategy. While DKF [28] addresses image processing delays, methods like PNG[30, 31, 32, 12] have been widely used to reduce overloads during the final interception phase. However, PNG is based on positional calculations, which may not be sufficient for non-cooperative target interception in real-world scenarios. Therefore, this paper designs a multicopter-friendly guidance law and controller capable of intercepting targets with low overload during the final interception phase, within the framework of existing IBVS sensors, using only the target angle information obtained from the monocular camera. By combining the above controller with our previously proposed DKF, we aim to achieve accurate interception of flight targets.
@@
In this paper, an IBVS controller is designed for the precise interception of non-cooperative flight targets. This task bears a similarity to that of natural bird predators (see Fig. 1). Specifically, the trajectory of a peregrine falcon hunting a mallard duck could be described by PNG [33], which mirrors the principles behind our approach. Moreover, a FOV holding controller mimics the bird’s neck rotation for prey tracking. Thus, the proposed controller can continuously and stably track the target and achieve precise interception, similar to a peregrine falcon’s hunting behavior. Overall, the contributions of this paper are summarized as follows:
@@
    This work introduces a PNG-based IBVS controller that utilizes only monocular camera measurements for target information, avoiding the assumptions of prior target knowledge [30, 32, 31] or the use of expensive and bulky LiDAR [12].
@@
    A controller for holding the field of view (FOV) is designed, incorporating PNG and multicopter dynamics to improve robustness in tracking targets under dynamic conditions.
@@
    The proposed method outperforms the classical PG [27] and state-of-the-art IBVS methods [28, 26] in interception accuracy for flight targets.
  4. 4)
    The system’s feasibility is validated through flight experiments, achieving a high interception success rate across diverse target motions and wind conditions.
@@ II-A Coordinate systems and relative motion models
There are four coordinate systems used, as shown in Fig. 2: the Earth-Fixed Coordinate System (EFCS): $o_{e}-x_{\mathrm{e}}y_{e}z_{\mathrm{e}}$, the Body Coordinate System (BCS): $o_{b}-x_{b}y_{b}z_{b}$, Camera Coordinate System (CCS): $o_{c}-x_{c}y_{c}z_{c}$ and Image pixel coordinate system (IPCS): $o_{i}-uv$.
@@ II-B Multicopter Dynamics Model
where ${}^{\mathrm{e}}\mathbf{p}_{\mathrm{M}}$ and ${}^{\mathrm{e}}\mathbf{v}_{\mathrm{M}}$ is the multicopter’s position and velocity in the EFCS, respectively. $\mathbf{g}$ denotes the acceleration due to gravity, typically represented as $\left[\begin{array}[]{ccc}0&0&g\end{array}\right]^{\mathrm{T}}$ with $g\approx$ 9.8 $\mathbf{m/s^{2}}$. ${}^{\mathrm{e}}\mathbf{f}\in\mathbf{R}^{3}$ is the controlled lift of the multicopter, directed opposite to the z-axis of the airframe. $\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\in\mathrm{SO}\left(3\right)$ represents the rotation from the BCS to the EFCS; ${}^{\mathrm{b}}\omega$ is the angular velocity in the BCS, and $\left[{}^{\mathrm{b}}\bm{\omega}\right]_{\times}$ is the cross-product matrix associated with ${}^{\mathrm{b}}\bm{\omega}$, where $\left[\cdot\right]_{\times}$ denotes the matrix such that $\left[\mathbf{x}\right]_{\times}\mathbf{y}=\mathbf{x}\times\mathbf{y}$ for any $\mathbf{x},\mathbf{y}\in\mathbf{R}^{3}$. $\mathbf{J}$ is the moment of inertia of the multicopter, $\mathbf{G}_{\mathrm{a}}$ represents the gyroscopic moment, and ${}^{\mathrm{b}}\mathbf{M}_{\mathrm{d}}$ denotes the aerodynamic moment associated with propeller steering.
@@ II-C Camera Image Model
where ${}^{\mathrm{c}}\tilde{\mathbf{v}}=\left[{}^{\mathrm{c}}\mathbf{v}^{\mathrm{T}}\,\,{}^{\mathrm{c}}\mathbf{\omega}^{\mathrm{T}}\right]^{\mathrm{T}}$ represents the instantaneous linear velocity angular and velocity of the camera. ${\bar{\mathbf{e}}}=\left[\bar{e}_{x}\,\,\bar{e}_{y}\right]^{\mathrm{T}}=\left[e_{x}/f_{oc}\,\,e_{y}/f_{oc}\right]^{\mathrm{T}}$ represents the normalized image error of features in IPCS, and $f_{\mathrm{oc}}$ is the focal length of the camera.
@@ II-D Framework overview
The system framework is shown in Fig. 3, the blue blocks indicate the core components of the proposed algorithm. First, the image is processed using a target detection algorithm [36], and corrections are estimated using the DKF [28]. The IBVS controller algorithm then calculates the desired lift $f_{\mathrm{d}}$ and angular velocity $\omega_{\mathrm{d}}$ for the multicopter flight controller. Finally, multicopter flight controller labeled green, such as PX4, generates PWM signals for the motors to execute the target interception.
@@ III-A Guidance Algorithm Design
$q_{z}=\arctan\left(\frac{n_{tx}}{n_{ty}}\right);q_{z}\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right).$
And normalized velocity direction vector $\mathbf{n}_{\mathbf{v}}=\left[n_{vx}\,\,n_{vy}\,\,n_{vz}\right]^{\mathrm{T}}$. Thus, $\sigma_{y}$ and $\sigma_{z}$ are the velocity angles in the vertical and horizontal planes which are denoted as:
@@
$\sigma_{z}=\arctan\left(\frac{n_{vx}}{n_{vy}}\right);\sigma_{z}\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right).$
@@
$\left\{\begin{array}[]{c}\sigma_{\mathrm{yd}}=K_{y}(q_{y(k)}-q_{y(k-1)})+\sigma_{y(k-1)}\\ \sigma_{\mathrm{zd}}=K_{z}(q_{z(k)}-q_{z(k-1)})+\sigma_{z(k-1)}\\ \end{array}\right.$ (9)
where $K_{y}$ and $K_{z}$ are proportionality coefficient for the vertical and horizontal planes. Layman et al. shows that a value of about 3 is appropriate when multirotor pursuit[37].
@@
where $\mathbf{n}_{\mathrm{vd}}=\left[\cos\sigma_{\mathrm{yd}}\sin\sigma_{\mathrm{zd}}\,\,\cos\sigma_{\mathrm{yd}}\cos\sigma_{\mathrm{zd}}\,\,\sin\sigma_{\mathrm{yd}}\right]^{\mathrm{T}}$. So the desired velocity can be get by designing the desired velocity magnitude.
@@ III-B Field of View Holding Controller
where $k_{\mathrm{p}}$ and $k_{\mathrm{d}}$ are small positive values. This controller improves the settling time and reduces the overshoot. Furthermore, it helps $e_{x}\rightarrow 0$ and is insensitive to pixel noise.
Changes in roll angle at greater distances from the target have less impact on the target within the FOV. Therefore, primary attention is given to changes in the pitch angle. To minimize the adverse effects of attitude changes, an acceleration controller is designed as follows:
@@
$\varDelta e_{y}\leqslant\frac{v_{0}\tan\left(\varDelta q_{\mathrm{y}}\right)}{\tan\left({\frac{1}{2}\alpha_{\mathrm{vfov}}}\right)}\leqslant\varepsilon$ (16)
@@ III-C Attitude Loop Controller
$\mathbf{a}_{\mathrm{d}}=\frac{\mathbf{v}_{\mathrm{d}}-\mathbf{v}_{\mathrm{now}}}{\mathrm{dt}}$ (17)
@@
$L_{1}=\mathrm{tr}\left(\mathbf{I}-\mathbf{R}_{\mathrm{d}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right).$ (20)
Given that $\left\|\mathbf{I}-\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right\|_{\mathrm{F}}=\sqrt{2\mathrm{tr}\left(\mathbf{I}-\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right)}$, $L_{1}=\left\|\mathbf{I}-\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}\right\|_{\mathrm{F}}^{2}/2\geq 0$, the derivative of the candidate function is:
$\dot{L_{1}}$ $=-\mathrm{tr}\left(\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}[^{\mathrm{b}}\bm{\omega}]_{\times}\right)$ (21)
$=\mathrm{vex}\left(\mathbf{R}_{\mathrm{d}}^{\mathrm{T}}\mathbf{R}_{\mathrm{b}}^{\mathrm{e}}-\mathbf{R}_{\mathrm{b}}^{\mathrm{eT}}\mathbf{R}_{\mathrm{d}}\right)^{\mathrm{T}}{{}^{\mathrm{b}}\bm{\omega}}.$
To make $\dot{L_{1}}\leqslant 0$ for stability, the pitch and roll angular velocity controllers are designed as:
@@
III-D Stability Analysis of the Proposed Controller
This subsection analyzes the closed-loop stability of the proposed image-based visual servo (IBVS) and proportional navigation guidance (PNG) fusion controller using Lyapunov’s direct method. In order to simplify the analysis while reflecting practical conditions, we make the following assumptions:
1) Initial Target Visibility. The target is initially within the camera’s field of view at a finite relative distance.
2) Target Maneuverability Limit. The target’s maximum velocity and acceleration are bounded such that LOS angle changes are dominated by the interceptor’s maneuvers.
Under these assumptions, we examine the tracking error dynamics and establish stability at the system level. We choose a Lyapunov candidate function as follows:
$L_{2}\;=\;\frac{1}{2}\Big{(}e_{x}^{2}\;+\;^{\mathrm{e}}p_{r}^{2}\Big{)}\!$ (24)
where ${}^{\mathrm{e}}p_{r}=|^{\mathrm{e}}\mathbf{p}_{\mathrm{r}}|$ is the relative distance to the target in the inertial frame. The derivative of Eq. (24) is
$\dot{L_{2}}\;=\;e_{x}\,\dot{e}_{x}\;+\;^{\mathrm{e}}p_{r}\,{}^{\mathrm{e}}\dot{p}_{r}\,.$ (25)
By construction, $L_{2}\geq 0$ for all defined error states, and $L_{2}=0$ if and only if $e_{x}={}^{\mathrm{e}}p_{r}=0$. According to Lyapunov’s second method, in order to make the system stable, $\dot{L_{2}}$ is expected to be negative definite.
For the image-plane error component, we approximate Eq. (12) as [26]
$\dot{e}_{x}=\frac{\bar{e}_{x}}{{}^{\mathrm{c}}\mathrm{p}_{tz}}{{}^{\mathrm{c}}v_{z}}-(1+{\bar{e}_{x}}^{2}){{}^{\mathrm{c}}\omega_{y}}.$ (26)
Considering ${{}^{\mathrm{c}}\omega_{y}}\approx{{}^{\mathrm{b}}w_{\psi}}$, substituting Eq. (13) into Eq. (26) yields a relationship between $e_{x}$ and $\dot{e}_{x}$ as
$\dot{e}_{x}=-\lambda{e}_{x}$ (27)
where
$\lambda=\frac{-\frac{{{}^{\mathrm{c}}v_{z}}}{{f_{oc}}^{\mathrm{c}}p_{tz}}+k_{p}\left(1+\bar{e}_{x}^{2}\right)}{1+k_{d}\left(1+\bar{e}_{x}^{2}\right)}.$
In order to make $\lambda>0$, we have make $k_{d}>0$ and $k_{p}>\frac{{}^{\mathrm{c}}v_{z}}{f_{oc}{{}^{\mathrm{c}}p_{tz}}\left(1+\bar{e}_{x}^{2}\right)}$. That is, $e_{x}$ and $\dot{e}_{x}$ always have opposite signs. Therefore, the first term in $\dot{L_{2}}$, $e_{x}\,\dot{e}_{x}$, is strictly negative.
Next, we examine the relative distance term. Neglecting the accelerations of both the interceptor and the target for analysis tractability, the LOS angular rate can be expressed by the relative motion geometry as
${}^{\mathrm{e}}p_{r}\,\dot{q}\;=\;v_{m}\sin\eta_{m}\;-\;v_{t}\sin\eta_{t}$ (28)
where $\eta_{m}=q-\sigma_{m}$ and $\eta_{t}=q-\sigma_{t}$ are the interceptor’s and target’s velocity leading angles, respectively. Differentiating the above relation and applying the PNG law $\dot{\sigma}_{m}=K\,\dot{q}$ for the interceptor yields
[…]
If the navigation constant $K$ is chosen to satisfy
$K\;>\;\frac{2\,\big{|}{}^{\mathrm{e}}\dot{p}_{r}\big{|}}{\,v_{m}\cos\eta_{m}\,}\,,$
then the coefficient $\big{(}K\,v_{m}\cos\eta_{m}+2\,{}^{\mathrm{e}}\dot{p}_{r}\big{)}$ becomes positive. In this case, $\ddot{q}$ is always of opposite sign to $\dot{q}$, which means the LOS rate $\dot{q}$ will monotonically decrease in magnitude and converge to zero. Therefore, $v_{m}\sin\eta_{m}-v_{t}\sin\eta_{t}$ of Eq. (28) converge to zero. Additionally, base on target maneuverability limit that $v_{t}<v_{m}$ and initial target visibility that $\cos\eta_{m}>0$, we can easily proved that [38]:
${}^{\mathrm{e}}\dot{p}_{r}\;=\;v_{t}\cos\eta_{t}\;-\;v_{m}\cos\eta_{m}\,<0.$ (30)
Since ${}^{\mathrm{e}}p_{r}>0$ by definition and ${}^{\mathrm{e}}\dot{p}_{r}<0$, the second term ${}^{\mathrm{e}}p_{r}\,{}^{\mathrm{e}}\dot{p}_{r}$ in $\dot{L_{2}}$ is also negative during the engagement.
In summary, since $L_{2}$ is positive definite and $\dot{L_{2}}<0$ for all nonzero errors, Lyapunov’s second method ensures the closed-loop error system asymptotically stable, guaranteeing the interceptor’s stable convergence to the target. However, the stability conclusions in this paper rely on the assumptions of limited target maneuverability and superior interceptor speed, with performance degradation observed under high-speed target escape conditions, as evidenced by experiments.